OpenStudy (anonymous):

how do you solve 12-4/5x-12=4 step by step please

OpenStudy (saifoo.khan):

\[12-\frac45x-12=4\]Like that?

jhonyy9 (jhonyy9):

12-12=0 -4x=20 x=20/-4 x=-5